If f(x,y,z)=xe^(2y)sin(9z), then the gradient is

∇f(x,y,z)=
1
Expert's answer
2020-10-19T15:43:45-0400

"\displaystyle

\nabla f(x, y, z) = \frac{\partial f(x, y, z)}{\partial x} \textbf{i} + \frac{\partial f(x, y, z)}{\partial y} \textbf{j} + \frac{\partial f(x, y, z)}{\partial z} \textbf{k} \\


\frac{\partial f(x, y, z)}{\partial x} = \frac{\partial (xe^{2y}\sin(9z))}{\partial x} = e^{2y}\sin(9z) \\

\frac{\partial f(x, y, z)}{\partial y} = \frac{\partial (xe^{2y}\sin(9z))}{\partial y} = 2xe^{2y}\sin(9z)\\

\frac{\partial f(x, y, z)}{\partial z} = \frac{\partial (xe^{2y}\sin(9z))}{\partial z} = 9xe^{2y}\cos(9z)\\

\therefore \nabla f(x, y, z) = e^{2y}\sin(9z) \textbf{i} + 2xe^{2y}\sin(9z)\textbf{j} + 9xe^{2y}\cos(9z)\textbf{k} \\

\therefore \nabla f(x, y, z) = e^{2y}(\sin(9z) \textbf{i} + 2x\sin(9z)\textbf{j} + 9x\cos(9z)\textbf{k})\\

\implies \nabla f(x, y, z) = e^{2y}(\sin(9z), 2x\sin(9z), 9x\cos(9z))"